OpenStudy (anonymous):

A man undertakes to do a piece of work in 150 days.he employs 200 men.he finds that only a quarter of work done in 50 days.how many additional men must be employed so that the work may be finished on time.( chain rule)

OpenStudy (anonymous):

let an additional x persons may be employed then by the prob (200+x)/200 =((1-1/4)/1/4)* 50/(150-100) or 1+x/200 =3/2 or x/200 =3/2 -1 or x=100

OpenStudy (anonymous):

here work and men are directly proportional whereas men and days are inversely proportional